hydroxylamine is a weak base with kb=7.96. determine the pH of 0.742 M hydroxylamine solution (to 2 decimals)
Answer #1

When weak base is tested for its pH

Equation is pOH = (Kb×C)1/2

pOH = ( 7.96×0.742) 1/2

pOH = 2.42

pH= 14-pOH = 14-2.42 = 11.58